OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8) I understand that the slope is 2 but i am not sure how to write out the final answer. I am stuck between these two answers y = 2(x + 2) y = 2(x – 2)

OpenStudy (anonymous):

y-0=2(x+2)

OpenStudy (anonymous):

So the answer is y=2(x+2)

OpenStudy (anonymous):

oh, that was simple thank you :)